OpenStudy (anonymous):

Can anyone help? what is the formula for amount of sodium that has been added in grams?

OpenStudy (unklerhaukus):

{mass of sodium} = {moles of sodium} times {molar mass of sodium} \[m_{Na} \small{[\text{g}]} =n_{Na} \small{[\text{mol}]}\times M_{Na}\small[\frac{\text{g}}{\text {mol}}]\] \[M_{Na}\small{[\frac{\text{g}}{\text {mol}}]}=23.99 \small[\frac{\text {g}}{ {\text{mol}}}]\]